# Segment as You Wish: Free-Form Language-Based Segmentation for Medical Images

### v1

@@ Abstract
Medical imaging is crucial for diagnosing a patient’s health condition, and accurate segmentation of these images is essential for isolating regions of interest to ensure precise diagnosis and treatment planning. Existing methods primarily rely on bounding boxes or point-based prompts, while few have explored text-related prompts, despite clinicians often describing their observations and instructions in natural language. To address this gap, we first propose a RAG-based free-form text prompt generator, that leverages the domain corpus to generate diverse and realistic descriptions. Then, we introduce FLanS, a novel medical image segmentation model that handles various free-form text prompts, including professional anatomy-informed queries, anatomy-agnostic position-driven queries, and anatomy-agnostic size-driven queries. Additionally, our model also incorporates a symmetry-aware canonicalization module to ensure consistent, accurate segmentations across varying scan orientations and reduce confusion between the anatomical position of an organ and its appearance in the scan. FLanS is trained on a large-scale dataset of over 100k medical images from 7 public datasets. Comprehensive experiments demonstrate the model’s superior language understanding and segmentation precision, along with a deep comprehension of the relationship between them, outperforming SOTA baselines on both in-domain and out-of-domain datasets.
1 Introduction
@@
Medical imaging is crucial in healthcare, providing clinicians with the ability to visualize and assess anatomical structures for both diagnosis and treatment. Organ segmentation is vital for numerous clinical applications, including surgical planning and disease progression monitoring [55, 14, 48]. However, accurately segmenting organs and tissues from these medical images, i.e., medical image segmentation (MIS), remains a significant challenge due to the variability in patient positioning, imaging techniques, and anatomical structures [42, 63]. Recent advancements in large foundation models, such as Segment Anything Model (SAM) [28] and MedSAM [62], have shown promise in achieving more accurate and faster MIS. These models often require the users to input a predefined category name, a box, or a point as a prompt. However, in real-world scenarios, clinicians often rely on natural language commands to interact with medical images, such as “Highlight the right kidney” or “Segment the largest organ”. An accurate segmentation model with flexible text comprehension capability is therefore essential for a wide range of clinical applications.
@@
The first challenge lies in the development of a segmentation model that can handle text prompts, offering greater flexibility and adaptability in real-world clinical environments. Unlike traditional models that rely on bounding boxes (Bboxes) or point prompts, this method should allow clinicians to use free-form natural language commands and streamline the diagnostic process by enabling intuitive, verbal interactions. For free-form text, we provide two conceptual definitions as follows: (1) Anatomy-Informed Segmentation, where the user has explicit knowledge of the organ or relevant pathology to be segmented; (2) Anatomy-Agnostic Segmentation, where the user lacks medical knowledge about a specific organ or CT scan and hence queries based on positional information, organ sizes or other visible characteristics. This scenario is more common for individuals such as students or patients without formal medical training. An exemplar illustration is shown below  [^1] All of the images in this paper are best view in color.:
{mdframed}
[backgroundcolor=gray!20, linecolor=gray!20, innertopmargin=5pt, innerbottommargin=5pt, innerleftmargin=10pt, innerrightmargin=10pt]
Example (1) : Anatomy-Informed Segmentation: An abdominal CT scan is recommended to evaluate for the presence of gallstones or any fluid accumulation around the gallbladder. — This professional diagnosis snippet indicates the most cared segmentation area is: Gallbladder.
Example (2): Anatomy-Agnostic Segmentation: I would love to have the leftmost organ segmented in this CT scan. — This description is agnostic to the medical name of the organ, but it indicates the segmentation target from positional semantics: Liver.
@@
To learn a free-form text-supportive MIS model, text prompt generation towards the groundtruth mask is a primary step. Instead of using labor-intensive manual labeling to match with the masks, we propose a retrieval augmented generation (RAG) fashion [30] method that automates text query generation using corpus embeddings collected from three resources (clinical expert records, non-expert queries, and synthetic queries). This approach guarantees that the generated query prompts capture various forms of language use across different demographic groups. Based on the text queries, we propose FLanS, a free-form language-based segmentation model that can accurately interpret and respond to free-form prompts either professional or straightforward, ensuring accurate segmentation across a variety of query scenarios.
Another challenge in text-based medical imaging segmentation arises from the variability in scan orientation. Factors such as patient positioning (e.g., supine vs. prone), different imaging planes (axial, coronal, sagittal), reconstruction algorithms and settings, and the use of portable imaging devices in emergency settings can cause organs to appear in unexpected locations or orientations. The scan orientations even differ between well-preprocessed datasets, such as AbdomenCT-1K [40] and BTCV [19], as shown in Fig. 1. This variability can confuse segmentation models, making it difficult to distinguish between the anatomical position of an organ and its appearance in a scan. For instance, the right kidney may appear either on the left or the right side of a rotated scan, leading to inaccurate segmentations. To address this challenge, we integrate the symmetry-aware canonicalization module as a crucial step in our model architecture [25, 41], which ensures the model produces consistent segmentations regardless of the scan’s orientation, enhancing its accuracy across diverse medical images [11, 60]. Additionally, incorporating symmetry improves sample efficiency and generalizability, which is well-suited for medical imaging tasks where labeled datasets are limited [53, 57, 72, 52].
Our key contributions in this paper are summarized as follows:
  • We employ RAG techniques for free-form text prompt generation for various anatomical structures containing diverse anatomy-informed and anatomy-agnostic queries. Stems from the vectorized embedding of clinical reports, produced query data employs the realistic tones and word usage.
  • We present a novel medical image segmentation model, FLanS, that exhibits a deep understanding of the relationship between text descriptions and medical images. It uniquely supports free-form text segmentation and employs a symmetry-aware canonicalization module to handle variability in scan orientation, as in Table. 1.
  • Our model training uses $\sim$100k medical images from 7 public datasets, covering 24 organs, along with diverse text prompts. This ensures the model generalizes across diverse anatomical structures and clinical scenarios and can be easily extended to new organs with upcoming datasets.
  • We demonstrate the FLanS’s effectiveness on both in-domain and out-of-domain datasets, and perform ablation studies to validate the contributions of each component in our model design.
@@
Table 1: FLanS uniquely supports all prompt types, including free-form text, and is symmetry-aware.
@@
SAM-U [12]
SAMed [68]
AutoSAM [24]
MedSAM [38]
MSA [62]
Universal  [33]
@@
Figure 1: Example CT images from different datasets show significant variations in orientation, which highlight the need for a symmetry-aware (equivariant) model to ensure consistent segmentation performance across diverse scan orientations.
2 Related Work
@@
Medical image segmentation (MIS) aims at accurately delineating anatomical structures in medical images. Traditionally, MIS methods tend to segment the correct regions from an image that accurately reflects the input query [2]. The researchers improve the performance of MIS methods by either optimizing segmentation network design for improving feature representations [8, 69, 7, 20], or improving optimization strategies, e.g., proposing better loss functions to address class imbalance or refining uncertain pixels from high-frequency regions to improve the segmentation quality [64, 49, 67]. However, they require a pre-known medical region from the user as an input for segmentation on where it is expected to be segmented and a precise match between the segment’s name and the labels used in the training set, restricting their flexibility in real-world application. Another category of methods are SAM-based approaches [28, 38, 71] that mainly rely on the Bboxes or points as prompts for segmentation. While such methods do not need strict labels, they neglect the descriptive understanding of the image, revealing a deficiency in performing arbitrary description-based segmentation, in comparison, our method handles well in Labels, free-form Text prompts without losing ability of Point and Bbox, as shown in the Table. 1.
@@
Text prompt segmentation, also referred to as expression segmentation [23], utilizes natural language expressions as input prompts for image segmentation tasks, moving beyond the traditional reliance on class label annotations [34]. Early research in this area employed CNNs and RNNs for visual and textual feature extraction, which were later combined through feature fusion for segmentation [31]. The success of attention mechanisms further inspired a new line of work [50, 66]. More recently, transformer-based architectures have improved segmentation performance by using either carefully designed encoder-based feature fusion modules [16, 65, 27] or decoder-based approaches [58, 35, 13]. Among these, [70] introduced a text-promptable mask decoder for efficient surgical instrument segmentation. However, there is no existing work that has focused on free-form language segmentation for diagnosis-related medical imaging tasks as introduced in this work.
@@
Equivariant neural networks ensure that their features maintain specific transformation characteristics when the input undergoes transformations, and they have achieved significant success in various image processing tasks [11, 60, 10, 4]. Recently, equivariant networks have also been applied to medical imaging tasks, including classification [61], segmentation [29, 15, 21], reconstruction [6], and registration [3]. Equivariance can be incorporated in different ways, such as through parameter sharing [17], canonicalization [25], and frame averaging [43]. In our work, since we leverage a pretrained segmentation network, we achieve equivariance/invariance through canonicalization [41], which, unlike other methods, does not impose architectural constraints on the prediction network. It uses a simple equivariant canonicalization network that transforms the input to a canonical form before feeding it to an unconstrained prediction network. By leveraging this technique, the performance and robustness of our model are greatly enhanced.
3 Methodology
@@
SAM [28] contains three main parts: (1) an image encoder that transforms images into image embeddings; (2) a prompt encoder that generates prompt embeddings; (3) a mask decoder that outputs the expected segmentation mask based on the image and prompt embeddings. Given a corresponding input medical image $x\in\mathcal{X}$ and a relevant prompt $p\in\mathcal{P}_{x}$. The image encoder embeds $x$ into $z_{x}$ that $z_{x}=\text{Encoder}^{\mathcal{X}}(x)$, similarly the prompt embedding $z_{p}=\text{Encoder}^{\mathcal{P}}(p)$. The mask decoder predicts the segmentation result (mask) by $\hat{m}_{x}^{p}=\text{Decoder}(z_{x},z_{p})$. While the SAM model provides $\text{Encoder}^{\mathcal{P}}$ for spatial prompts (e.g. Bbox or point), the integration of text-based prompts has been less explored. In text-based medical images segmentation, natural language prompts require specialized learning to effectively capture clinical terminology and segmentation intent.
3.1 The Retrieval Augmented Query Generator
@@
To equip a MIS model $\mathcal{M}$ with language comprehension abilities, it is essential to prepare a suitable natural language query  [^2] Throughout the paper, we use the terms “query” and “prompt” interchangeably. corpus $\mathcal{C}$ in correspondence with the target
Figure 2: The RAG Free-form Query Generator. The domain corpus, from the EMRs embedding, completes the retrieval augmentation and enhances the LLMs with the clinical way of query.
@@
For $\mathcal{S}_{1}=\textit{Domain Expert}$, we collected over 7,000 reports written by doctors and identified 4,990 clinical diagnosis records that are relevant to 24 labeled organs for this study. After de-identification, we embed such Electronic Medical Records (EMRs) into semantic vector space through Med-BERT [45], which outperforms the general language embedding models such as Bert or GPTs in the bioinformatics context understandings. Then, we built a retrieval augmented generation fashion generator agent $\mathbf{G}$, as shown in Fig. 2, provided with medical domain corpus and practitioner’s language usage preference. It retains the original LLM’s natural language ability such as sentences extension and rephrasing. Finally, we construct a query prompt template: “System: You are an agent able to query for segmenting label {Liver} in this {CT} scan. Please write the query sentence and output it.” Given a label $l_{i}$ = Liver, where $l\in\mathcal{L}$ regarding an arbitrary organ label with CT modality, the $\mathbf{G}$ produces a free-form query $q_{l}^{i}$, this query is taken as prompt in the later text-aware segment model training. E.g., “(1) Examine this CT scan to determine the extent of hepatic damage present. (2) As the symptoms suggest cirrhosis, we should analyze the related part in this CT scan for any signs of the disease. These retrieved augmented results show that the interested organ may not always be explicitly mentioned, but can be inferred based on terms like ‘cirrhosis’ and ‘hepatic’, which are all liver-specific illnesses in clinical practice.
@@
Figure 3: Spatial features extracted from the Bboxes of ground truth masks are processed by the RAG query generator $\mathbf{G}$ to produce anatomy-agnostic queries.
3.2 Free-Form Language Segmentation for Medical Images

### v2

# FLanS: A Foundation Model for Free-Form Language-based Segmentation in Medical Images

@@ Abstract.
Medical imaging is crucial for diagnosing a patient’s health condition, and accurate segmentation of these images is essential for isolating regions of interest to ensure precise diagnosis and treatment planning. Existing methods primarily rely on bounding boxes or point-based prompts, while few have explored text-related prompts, despite clinicians often describing their observations and instructions in natural language. To address this gap, we first propose a RAG-based free-form text prompt generator, that leverages the domain corpus to generate diverse and realistic descriptions. Then, we introduce FLanS, a novel medical image segmentation model that handles various free-form text prompts, including professional anatomy-informed queries, anatomy-agnostic position-driven queries, and anatomy-agnostic size-driven queries. Additionally, our model also incorporates a symmetry-aware canonicalization module to ensure consistent, accurate segmentations across varying scan orientations and reduce confusion between the anatomical position of an organ and its appearance in the scan. FLanS is trained on a large-scale dataset of over 100k medical images from 7 public datasets. Comprehensive experiments demonstrate the model’s superior language understanding and segmentation precision, along with a deep comprehension of the relationship between them, outperforming SOTA baselines on both in-domain and out-of-domain datasets. For code and models, please find them in the released project page by clicking here.
Foundation Model, Medical Image, Language Based Segmentation
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736963isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Medical image segmentationccs: Computing methodologies Neural networksccs: Computing methodologies Language modelsccs: Information systems Retrieval augmented generation
1. Introduction
@@
Medical imaging is crucial in healthcare, providing clinicians with the ability to visualize and assess anatomical structures for both diagnosis and treatment. Organ segmentation is vital for numerous clinical applications, including surgical planning and disease progression monitoring (Wang et al., 2022a; Du et al., 2020; Shamshad et al., 2023). However, accurately segmenting organs and tissues from these medical images, i.e., medical image segmentation (MIS), remains a significant challenge due to the variability in patient positioning, imaging techniques, and anatomical structures (Pham et al., 2000; Xiao and Sun, 2021). Recent advancements in large foundation models, such as Segment Anything Model (SAM) (Kirillov et al., 2023) and MedSAM (Wu et al., 2023), have shown promise in achieving more accurate and faster MIS. These models often require the users to input a predefined category name, a box, or a point as a prompt. However, in real-world scenarios, clinicians often rely on natural language commands to interact with medical images, such as “Highlight the right kidney” or “Segment the largest organ”. An accurate segmentation model with flexible text comprehension capability is therefore essential for a wide range of clinical applications.
Figure 1. Examples of two types of language-prompted segmentation. (1) A professional diagnosis snippet from EMR uses medically relevant information to indicate that the most important segmentation area is the Gallbladder. (2) This description is agnostic to the medical name of the organ but identifies the segmentation target based on positional semantics: the Liver.
@@
The first challenge lies in the development of a segmentation model that can handle text prompts, offering greater flexibility and adaptability in real-world clinical environments. Unlike traditional models that rely on bounding boxes (Bboxes) or point prompts, this method should allow clinicians to use free-form natural language commands and streamline the diagnostic process by enabling intuitive, verbal interactions. For free-form text, we provide two conceptual definitions as follows: (1) Anatomy-Informed Segmentation, where the user has explicit knowledge of the organ or relevant pathology to be segmented; (2) Anatomy-Agnostic Segmentation, where the user lacks medical knowledge about a specific organ or CT scan and hence queries based on positional information, organ sizes or other visible characteristics. This scenario is more common for education purposes that individuals (like students) without formal medical training. An exemplar illustration is shown in Fig. $1$All of the images in this paper are best viewed in color.
@@
To learn a free-form text-supportive MIS model, text prompt generation towards the groundtruth mask is a primary step. Instead of using labor-intensive manual labeling to match with the masks, we propose a retrieval augmented generation (RAG) (Lewis et al., 2020) fashion method that automates text query generation using corpus embeddings collected from three resources (clinical expert records, non-expert queries, and synthetic queries). This approach guarantees that the generated query prompts capture various forms of language use across different demographic groups. Based on the text queries, we propose FLanS, a free-form language-based segmentation model that can accurately interpret and respond to free-form prompts either professional or straightforward, ensuring accurate segmentation across a variety of query scenarios.
@@
Table 1. FLanS uniquely supports all prompt types, including free-form text, and is symmetry-aware.
@@
SAM-U (Deng et al., 2023)
SAMed (Zhang and Liu, 2023)
AutoSAM (Hu et al., 2023)
MedSAM (Ma et al., 2024)
MSA (Wu et al., 2023)
Universal  (Liu et al., 2023b)
@@
Figure 2. Example CT images from different datasets show variations in orientation, which highlight the need for a symmetry-aware (equivariant) model to ensure consistent segmentation performance across diverse scan orientations.
Another challenge in text-based medical imaging segmentation arises from the variability in scan orientation. Factors such as patient positioning (e.g., supine vs. prone), different imaging planes (axial, coronal, sagittal), reconstruction algorithms and settings, and the use of portable imaging devices in emergency settings can cause organs to appear in unexpected locations or orientations. The scan orientations even differ between well-preprocessed datasets, such as AbdomenCT-1K (Ma et al., 2022) and BTCV (Gibson et al., 2018), as shown in Fig. 2. This variability can confuse segmentation models, making it difficult to distinguish between the anatomical position of an organ and its appearance in a scan. For instance, the right kidney may appear either on the left or the right side of a rotated scan, leading to inaccurate segmentations. To address this challenge, we integrate the symmetry-aware canonicalization module as a crucial step in our model architecture (Kaba et al., 2022; Mondal et al., 2023), which ensures the model produces consistent segmentations regardless of the scan’s orientation, enhancing its accuracy across diverse medical images (Cohen and Welling, 2016; Weiler and Cesa, 2019). Additionally, incorporating symmetry improves sample efficiency and generalizability, which is well-suited for medical imaging tasks where labeled datasets are limited (Wang et al., 2022c, 2021; Zhu et al., 2022; Thomas et al., 2018).
Our key contributions in this paper are summarized as follows:
  • We employ RAG techniques to generate free-form text prompts for various anatomical structures, incorporating both anatomy-informed and anatomy-agnostic queries. Derived from the vectorized embeddings of clinical reports, the generated query data adopts realistic tones and natural word usage.
  • We introduce FLanS, a novel medical image segmentation model that demonstrates a deep understanding of the relationship between text descriptions and medical images. It uniquely supports free-form text segmentation and employs a symmetry-aware canonicalization module to handle variability in scan orientation, as illustrated in Table. 1.
  • We demonstrate the effectiveness of FLanS across diverse anatomical structures and clinical scenarios using both in-domain and out-of-domain datasets. Additionally, we conduct comprehensive ablation studies to validate the contribution of each component in our model design.
2. Related Work
@@
Medical image segmentation (MIS) aims at accurately delineating anatomical structures in medical images. Traditionally, MIS methods tend to segment the correct regions from an image that accurately reflects the input query (Azad et al., 2024). The researchers improve the performance of MIS methods by either optimizing segmentation network design for improving feature representations (Chen et al., 2018b; Zhao et al., 2017; Chen et al., 2018a; Gu et al., 2020), or improving optimization strategies, e.g., proposing better loss functions to address class imbalance or refining uncertain pixels from high-frequency regions to improve the segmentation quality (Xue et al., 2020; Shi et al., 2021; You et al., 2022). However, they require a pre-known medical region from the user as an input for segmentation on where it is expected to be segmented and a precise match between the segment’s name and the labels used in the training set, restricting their flexibility in real-world application. Another category of methods are SAM-based approaches (Kirillov et al., 2023; Ma et al., 2024; Zhu et al., 2024) that mainly rely on the Bboxes or points as prompts for segmentation. While such methods do not need strict labels, they neglect the descriptive understanding of the image, revealing a deficiency in performing arbitrary description-based segmentation. In comparison, our method handles well in Labels, free-form Text prompts without losing the ability of Point and Bbox, as shown in the Table. 1.
@@
Text prompt segmentation, also referred to as expression segmentation (Hu et al., 2016), utilizes natural language expressions as input prompts for image segmentation tasks (Liu et al., 2023a), moving beyond the traditional reliance on class label annotations (Liu et al., 2021), such as nn-Unet (Isensee et al., 2018), and Swin-unet (Cao et al., 2022). Early research in this area employed CNNs and RNNs for visual and textual feature extraction, which were later combined through feature fusion for segmentation (Li et al., 2018). The success of attention mechanisms further inspired a new line of work (Shi et al., 2018; Ye et al., 2019). More recently, transformer-based architectures have improved segmentation performance by using either carefully designed encoder-based feature fusion modules (Feng et al., 2021; Yang et al., 2022; Kim et al., 2022) or decoder-based approaches (Wang et al., 2022b; Lüddecke and Ecker, 2022; Ding et al., 2021). Among these, (Zhou et al., 2023) introduced a text-prompt mask decoder for efficient surgical instrument segmentation and  (Zhao et al., 2024) introduced BiomedParse to perform universal segmentation tasks in one model by text guidance. However, there are few existing works that have focused on free-form language segmentations, and the existing work (Zhao et al., 2024) spends much labor on text annotations, we provide a more efficient domain-specific solution by using retrieval augmented generation in this work.
@@
Equivariant neural networks ensure that their features maintain specific transformation characteristics when the input undergoes transformations, and they have achieved significant success in various image processing tasks (Cohen and Welling, 2016; Weiler and Cesa, 2019; Cohen et al., 2019; Bronstein et al., 2021). Recently, equivariant networks have also been applied to medical imaging tasks, including classification (Winkels and Cohen, 2018), segmentation (Kuipers and Bekkers, 2023; Elaldi et al., 2024; He et al., 2021), reconstruction (Chen et al., 2021), and registration (Billot et al., 2024). Equivariance can be incorporated in different ways, such as through parameter sharing (Finzi et al., 2021), canonicalization (Kaba et al., 2022), and frame averaging (Puny et al., 2021). In our work, since we leverage a pretrained segmentation network, we achieve equivariance/invariance through canonicalization (Mondal et al., 2023), which, unlike other methods, does not impose architectural constraints on the prediction network. This paper is the first to leverage the Equivariant methods to align the orientation feature in the image with the text description and achieve good medical image segmentation performance by resolving the ambiguities.
3. Methodology
@@
SAM (Kirillov et al., 2023) contains three main parts: (1) an image encoder that transforms images into image embeddings; (2) a prompt encoder that generates prompt embeddings; (3) a mask decoder that outputs the expected segmentation mask based on the image and prompt embeddings. Given a corresponding input medical image $x\in\mathcal{X}$ and a relevant prompt $p\in\mathcal{P}_{x}$. The image encoder embeds $x$ into $z_{x}$ that $z_{x}=\text{Encoder}^{\mathcal{X}}(x)$, similarly the prompt embedding $z_{p}=\text{Encoder}^{\mathcal{P}}(p)$. The mask decoder predicts the segmentation result (mask) by $\hat{m}_{x}^{p}=\text{Decoder}(z_{x},z_{p})$. While the SAM model provides $\text{Encoder}^{\mathcal{P}}$ for spatial prompts (e.g. Bbox or point), the integration of text-based prompts has been less explored. In text-based medical images segmentation, natural language prompts require specialized learning to effectively capture clinical terminology and segmentation intent.
Figure 3. The RAG Free-form Query Generator. The domain corpus, from the EMRs embedding, completes the retrieval augmentation and enhances the LLMs with the clinical way of query.
3.1. The Retrieval Augmented Query Generator
Figure 4. Spatial features extracted from the Bboxes of ground truth masks are processed by the RAG query generator $\mathbf{G}$ to produce anatomy-agnostic queries. For sample $x$, spatial info. is extracted for mask $m_{x}^{(i)}$ using Bboxes, classifying to spatial categories: largest, smallest, left-most, right-most, etc. The RAG generator $\mathbf{G}$ then converts categories into anatomy-agnostic text queries to augment $\mathcal{P}_{x}$.
@@ 3.1. The Retrieval Augmented Query Generator
For $\mathcal{S}_{1}=\textit{Domain Expert}$, we collected over 7,000 reports written by doctors and identified 4,990 clinical diagnosis records that are relevant to 24 labeled organs for this study. After de-identification, we embed such Electronic Medical Records (EMRs) into semantic vector space through Med-BERT (Rasmy et al., 2021), which outperforms the general language embedding models such as Bert or GPTs in the bioinformatics context understandings. Then, we built a retrieval augmented generation fashion generator agent $\mathbf{G}$, as shown in Fig. 3, provided with medical domain corpus and practitioner’s language usage preference. It retains the original LLM’s natural language ability such as sentences extension and rephrasing. Finally, we construct a query prompt template: “System: You are an agent able to query for segmenting label {Liver} in this {CT} scan. Please write the query sentence and output it.” Given a label $l_{i}$ = Liver, where $l\in\mathcal{L}$ regarding an arbitrary organ label with CT modality, the $\mathbf{G}$ produces a free-form query $q_{l}^{i}$, this query is taken as prompt in the later text-aware segment model training. E.g., “(1) Examine this CT scan to determine the extent of hepatic damage present. (2) As the symptoms suggest cirrhosis, we should analyze the related part in this CT scan for any signs of the disease. These retrieved augmented results show that the interested organ may not always be explicitly mentioned, but can be inferred based on terms like ‘cirrhosis’ and ‘hepatic’, which are all liver-specific illnesses in clinical practice.
@@
Figure 5. The architecture of our proposed model FLanS. First, given a set of free-form text prompts $p_{1}...,p_{n}$, the text encoder gets the text embedding, and then passes through the learned Intention Head Layer that maps the embedding to a space with explicit intention probabilities, which is useful for the FLanS model weight updating as in Eq. 1. Second, we have trained a Canonicalized Network that transforms any medical image with arbitrary orientation into a canonicalization space, making sure the encoded image aligns with the standard clinical practice to avoid ambiguity. Third, the encoded prompts (either spatial info such as Point, Bbox, or Free-form text data), together with the encoded image, will be processed with mask decoder and output the expected masks.
3.2. Free-Form Language Segmentation